08:42 Chang of running conditions. Because no carbon events were seen we're going back to 3He running only. Changing target to 3He. Will restart with radiator in.
08:52 Target in requested 30uA for radiator in running.
09:55 We're still getting transient beeps from the HRS magnets, but it seems better than in the afternoon. They must be late risers.
09:58 J. Segal wants to clear some latches on HRS-L Q3
10:48 Paged J. Segal to clear his latches. Cleared without incident.
11:44 HRS-R Q2 drifted low. Stopped run # 1390 and magnet recovered on its own.
11:49 Just started run 1391 and the magnet drifted again. Paged J. Larose for some advice. He suggested making a slight change, allowing some time to settle and then resetting to the desired setting. We wuill try that and see how it goes.
11:54 Restarting with run #1392. Things seem to have settled down a bit.
12:51 Changing to radiator out
12:53 Resumed running with radiator out, 3He target, 50uA.
12:57 Accidently knocked the HRS-L dipole out of regulation trying to get the HRS-L Q3 to settle down! Darn it @! On the up-side, Q3 did settle down.
13:01 MCC called to let us know that they had a spectrometer trip off, and they need to reset it.
13:04 The HRS-L is back up and running, mercifully we didn't loose too much since MCC had their problem at about the same time we did.
13:05 MCC reports that the spectrometer magnet has a water leak and it will take at least 30 minutes to recover.
13:42 MCC called to let us know that they are ready to send us beam. We requested 50uA. Then there was an FSD.
14:12 The beam seems to be a little more unstable since the spectrometer trip.
14:36 DAQ crashed: TS0 crashed. Rebooting. Effected run #1394
14:38 Back to running. Run #1395, Radiator out.
15:30 Changed to radiator in 6%. Requested 30uA
16:00 Finished shift running radiator in, 3He target, 30uA.